OpenStudy (fanduekisses):

Convert from radical to exponential form, is this correct?(wait till I finish writing the formulas please.)

OpenStudy (fanduekisses):

\[\sqrt[5]{32d ^{3}e ^{5}f} = 2^{\frac{ 1 }{5 }}d ^{\frac{ 3 }{ 5 }} f ^{\frac{ 1 }{ 5 }}\]

OpenStudy (anonymous):

The variables are okay.. But 32 is 2^5....the 5th root of 2^5 is just 2...so you don't need an exponent on your 2 in the answer.

OpenStudy (anonymous):

Think of it as (2^5)^1/5 which is 2^(5*1/5) which is 2^1 or 2

OpenStudy (anonymous):

@BangkokGarrett Where is the e?

OpenStudy (anonymous):

ooops you forgot the e. there should be an e in the answer.

OpenStudy (fanduekisses):

oops yeah I forgot the e

OpenStudy (anonymous):

(e^5)^1/5 = e^(5*1/5) = e^1 = e

OpenStudy (fanduekisses):

I'm confused

OpenStudy (fanduekisses):

help